OpenStudy (anonymous):

help please Y = x + 3 , in point-slope form

OpenStudy (anonymous):

what do you need to know? the slope is 1/1 and the y-intercept is positive 3.

OpenStudy (anonymous):

can yah write out the steps please thank you all

OpenStudy (anonymous):

i needed to write it out in point slope form

OpenStudy (anonymous):

ill give a medal btw

OpenStudy (anonymous):

y-y1=m(x-x1). in this equation y=3. x=3 so its y-3=1/1(x-3)

OpenStudy (anonymous):

i read that before it was a little hard to understand :?

OpenStudy (anonymous):

you figure out x by graphing it until it crosses the x axis

OpenStudy (anonymous):

does mine help?

OpenStudy (anonymous):

thx

OpenStudy (anonymous):

yup